# Readout electronics for low occupancy High-Pressure Gas TPCs

### v1

@@ 2.2 Firmware design
The PAT card’s main processing unit is a Kintex Ultrascale FPGA package. This device was selected as it provides the required number of serial transceivers to accommodate both readout schemes for ND-GAr whilst satisfying the remaining requirements. The power consumption (based on the AMD power estimate tool) is within the nominal limits for operation in gases at atmospheric pressure or above when actively cooled. Finally, the availability of this package at low cost through CERN and the extensive experience of using it that has been built up at CERN make this FPGA a favourable choice given the cost driven nature of the design.
In order to take advantage of the FPGA, dedicated firmware had to be developed that allowed data to be reliably and efficiently streamed. Here, we discuss the firmware solution for direct powering, as this was the firmware that was eventually tested with the TOAD test beam. The design has two primary functions. First, the data flow, i.e. how FEC data are reliably received into the software. The second is Control Configuration and Monitoring (CCM) operations. A schematic representation of the firmware design can be seen in figure 4.
@@ 2.2.1 Data flow
The data stream is then decoded using dedicated decoder firmware based on a Finite State Machine (FSM) that can handle all the different operational modes of the SAMPA. The extracted payload along with some of the header information (channel number, timestamp, and number of clusters), is converted into a stream that is subsequently sent through a RR based buffering system. The output Ethernet sample contains additionally a single header with some of the header information received from the SAMPA header, as well as the DUNE timestamp for data alignment.
@@
The second option interfaces the SiTCP suite from BeeBeans [14] with IPbus. SiTCP is a lightweight interface that provides streaming capabilities over Transmission Control Protocol (TCP). We developed an interface between the standard pull-request based IPbus firmware and the SiTCP IP core, thereby providing a TCP based interface for the IPbus suite. While SiTCP includes its own UDP based interface for CCM, it does not have enough address space for our application.
The third option was used for the power over serial powering mode investigations. The number of serial links on the PAT only allows for a single hardware interface between the in-pressure and out-of-pressure boards for both data and CCM and hence the SiTCP solution can be used in the out-of-pressure electronics. The link between the in-pressure and out-of-pressure electronics still cannot use the SiTCP solution, so it is implemented as an Aurora based stream interface. In this operation mode, IPbus is not used at all and the received data are converted into a TCP stream interface using SiTCP.
@@ 2.2.2 Control, Configuration and Monitoring
The firmware includes many registers containing information allowing full telemetry of the hardware from software. The CCM aspect of the firmware’s job is to monitor and set these registers. The design of the PAT cards allows the FPGA to also monitor and change the behaviour of several other components on the board using standardised slow control interfaces (e.g. Inter-Integrated Circuit (I2C), Serial Peripheral Interface (SPI), Power-Management Bus (PMbus)). The list of components can be found in table 5 along with a short explanation of their function.
@@
In addition to the external hardware components, the CCM part of the firmware implements a variety of standard IPbus registers and memories dedicated to monitoring the data flow or configuring operational modes. A list of these registers can be seen in table 6.
@@
| Sync. control | Send synch signal to all FEC sampling windows | 1/FEC group |
@@ 2.3 Software development
The data flow application is based on the DUNE DAQ software suite. For CCM, an IPbus based framework has been developed allowing full access to all the in-pressure electronic’s configuration registers. Finally, for timing, the software is based on the framework from the DUNE timing system which we have wrapped in the bespoke PAT CCM framework since the timing system uses IPbus for configuration as well.
@@ 2.3.1 DUNE DAQ integration
Once the data has been streamed out, it is received by the DUNE DAQ software. While this software already existed prior to this work, significant effort was required to integrate TOAD’s DAQ into the larger DUNE software stack. A schematic of the general data flow within DUNE DAQ is shown in figure 6. The integration of the TOAD readout into this framework involved adding the Ethernet packet structure described in section 2.2.1 into the upstream DAQ. Additionally, the zero-suppression functionality of the FECs allows for the PAT boards to only output data above the threshold we wish to be considered for offline processing. This means that TOAD also demonstrated that the DUNE DAQ software framework worked in untriggered continuous readout mode.
A further difference between TOAD and the other systems using DUNE DAQ is that the output stream data format allows each hit that is streamed out to contain a variable number of ADC samples, while all other systems had fixed numbers of samples per transmitted data packet. This implementation expands the capabilities of the DUNE DAQ software and allows it to handle variable length data packets[12]. We carried out a test of this implementation and verified that it worked using real PAT boards and data from simulated muon tracks in an HPgTPC.
@@ 2.3.2 Control, Configuration and Monitoring software suite
The data from TOAD is received by standard DUNE software [12]. However, the CCM software suite is a new scalable and modular software framework. The framework has two key base classes to talk to the two classes of objects on the PAT boards: firmware-implemented registers and external devices. The framework is summarised in figure 7. Each base class contains general methods needed by all registers of that type. Communication to the firmware registers is via IPBus, so the base class contains IPBus specific methods, while the inherited classes for each specific register contain their own specific bit maps that relate their 32-bits to specific functions. This scheme allows scalability to all electronic devices that use IPBus and similar firmware implementation.
@@
The CCM suite also provides various low-level scripts that act as a Command Line Interface (CLI) for the user to manage the hardware. By combining multiple low-level scripts, one can create high-level scripts that are tailored towards specific use cases. Examples include the noise measurements outlined in section 4.1 and continuous temperature and power monitoring used in section 4.2.
3 Test environment
3.1 TOAD
The TOAD detector design and operation largely follow that of the UK high-pressure platform in [9]. Our amplification wire chamber comes from the ALICE [16, 6] at the CERN LHC, which studies quark-gluon plasma in high-energy Pb-Pb collisions. The ALICE TPC is the main component of the experiment for tracking and particle identification, requiring the ability to track particle densities of up to per unit of rapidity at , with a maximum trigger rate of [8]. The TPC’s shape is that of a hollow cylinder with a length of and radius of , leaving space in the centre for ALICE’s cylindrical inner tracking system with a radius of . The readout plane at each endplate is segmented into 18 trapezoidal sectors, with an Inner Read-Out Chamber (IROC) and an OROC per sector. Each ROC is a MWPC. Given the increased Pb-Pb luminosity delivered to the LHC in Run 3, corresponding to a collision rate of , the ALICE TPC required an upgrade to avoid pile-up [7]. Hence, the ROCs can be repurposed and an OROC is used as the readout chamber for the TOAD TPC.
3.2 Pressure vessel
TOAD utilises the same pressure vessel described in [9, 17], which allows the testing of the readout electronics with an ALICE OROC at pressures up to (). The vessel has an inner diameter of and an outer diameter of , made from Stainless Steel. The vessel has domed ends, giving a total volume of . The vessel can be opened by removing one of the domed faces. The vessel door is secured to the vessel body via a DN 1500 flange, where the flange on the vessel body has two grooves for O-rings: a diameter circular viton inner ring, and a rectangular silicon outer ring. The inner ring ensures chamber pressure tightness in vacuum conditions and the outer ring ensures it in high-pressure conditions. A detailed schematic of the vessel and its dimensions is shown in figure 8 and an image of the vessel in the test environment is shown in figure 9. There are five DN200 flanges on the vessel door, which are not used. One DN200 flange on the back of the vessel is used to mount a camera via a thick quartz optical window. There are various KF25 and KF40 flanges on the vessel, which are used for the High Voltage (HV), gas, vacuum and readout-electronics systems feedthroughs. The gas and evacuation system is connected to the vessel so that the vessel can be evacuated to . This is achieved by first evacuating with a roughing pump, and once the vessel is below , evacuation with a turbo pump through a Carten solenoid valve. The pump system used is an Edwards Turbo Cart which consists of a backing pump, turbo pump and controller. The pressure inside the vessel is monitored using two pressure sensors - an Inficon PGC550 combined capacitance-Pirani vacuum gauge for below-atmosphere pressure readings and a Wika A-10 digital pressure gauge for above-atmosphere pressure readings ( to ). These are attached directly to the vessel via feedthroughs and readings are stored regularly through the slow control system. The gas system allows the vessel to be filled with up to of gas mixtures, with the gas mix controlled by filling to a given partial pressure of gas from up to four different gas cylinders connected to the system. For the run described here, we were able to mix the following three input gases: premixed Ar-CH4 with molar ratio 92:8 for Ar:CH4, pure Ar, and pure N2.
The TPC consists of a HV drift cathode, which is a steel mesh ring of radius , made with wires of diameter with 25 lines per inch. There are three HV feedthroughs on the vessel, which are used to apply voltages to the cathode, anode and gating grid. Other feedthroughs are also used to ensure the equipment inside the vessel is all safely grounded. The cathode and OROC anode have been tested with voltages up to 16 kV and 3 kV respectively, with the maximum achievable voltage varying with the gas mixture used [9]. The field cage consists of eight copper rings of inner diameter , connected by resistors, with resistors of this resistance also connecting the first ring to the drift cathode and the last ring to ground [17]. The results shown in this paper use various configurations of HV, gas mixtures and pressures. These are summarised in table 7.
| 1 | Air | 1.0 | 0 | 0 | 0 |
| 2 | Ar:CH4, 96:4 | 4.5 | 0 | 0 | 0 |
| 3 | Ar:CH4, 98:2 | 3.0 | 1500.0 | -100.0 | 0 |
| 4 | Ar:CH4, 98:2 | 3.0 | 1500.0 | -100.0 |
In addition to the connections described above, the readout electronics were powered and data streamed through KF40 feedthroughs on the vessel door. One feedthrough was used to provide power to all three PAT boards. Each board then communicates through an individual RJ45 feedthrough connected directly to a COTS network switch which is, in turn, connected to two DELL PowerEdge servers: one running CCM software, and the other DAQ software – both of which are outlined in detail in section 2.3. Other feedthroughs on the vessel body are used for safety burst discs – a burst disc backed by a pressure relief valve and an independent burst disc direct to atmosphere – to ensure the vessel pressure does not exceed its rated pressure.
3.2.1 ALICE Multi-Wire Proportional Chambers
The pre-upgrade ALICE ROCs employ a typical wire plane schematic: an anode-wire grid above the pad plane, a cathode-wire grid and a gating grid to separate the drift volume from the readout. The pad plane and the three wire planes are all spaced apart. All wires run in the azimuthal direction (approximately vertically in the orientation we have the chamber) and have different separations in the various planes. The length of the shorter and longer parallel sides of the OROC are and , respectively. The perpendicular length of the OROC is . The gating grid wires have a pitch (spacing) of and are made of copper beryllium with a diameter. The cathode wires have the same properties but with a pitch of and are staggered to the gating grid wires. Likewise, the anode wires are staggered compared to the cathode and gating grid wires, with a pitch of . These anode wires, in contrast to the other two planes, must withstand a HV and are made from gold-plated tungsten with a diameter. The edge wires of the anode (first two and last two) have a thicker diameter of . A gating grid wire voltage of or is set to alternate wires. In this way, drift electrons or ions cannot pass the gating grid region. The cathode wires are held at ground. However, as the OROC used in TOAD was part of the ALICE ‘test’ ROCs, rather than a production ROC, the ability to change the bias at the cathode wires is a feature. The wires can be seen on the OROC in figure 10(a). The OROC has previously been successfully tested up to anode voltages of in Ar-CO2 and Ar-CH4 mixtures with pressures above [9]. The pad plane consists of 9984 pads in total, of two sizes, which are summarised in table 8. Here, “inner” refers to the rows closer to the shorter parallel side, whereas “outer” refers to rows closer to the longer side of the OROC. The difference in pad dimensions is due to the trapezoidal shape of the ROC. Groups of 21 or 22 pads are read out via a socket on the back of the OROC and the sockets are grouped in threes. As shown in figure 2, a TOAD FEC has three connectors, meaning that it can plug into three sockets and hence reads out 64 pads. The FECs and PAT boards are shown connected to the OROC in figure 10(b).
| Pad Dimensions (mm2) | Pad Rows | Number of Pads | |
| Outer Pads | 6 15 | 32 | 4032 |
| Inner Pads | 6 10 | 64 | 5952 |
| Total | - | 96 | 9984 |
In the TPC, ionisation electrons are produced and then drift towards the ROC due to the presence of an electric field, from the HV cathode electrode and the gating grid. When the electrons reach the ROC, they undergo avalanches at the anode wires and therefore, charge is induced onto the pads. The spatial position of the ionisation electrons in the ROC plane is determined from the pad signals. The perpendicular position, in the drift direction, is calculated from the comparative time of the signal on the ROC and a time, , which would come from beamline instrumentation triggers. The studies outlined in this paper are noise studies and, therefore, particle tracking is not employed.
@@ 4.1 Noise characterisation
Understanding the noise profile is crucial in verifying that we have a sufficient signal-to-noise ratio to identify the signals we are looking for. We characterise noise as the Root Mean Square (RMS) error of the waveform in the absence of any signal.
@@
We compute noise by taking the RMS of a waveform, which corresponds to 500 samples at , for each channel in each FEC. Figure 11 shows the spatial distribution of the RMS noise of all FECs in the OROC, measured under pressure. Channels with RMS noise less than are coloured grey, while channels with noise above are shown with the highest colour in the scale. The blank regions are caused by FECs that did not respond to any I2C communications.
@@
Figure 12 plots the distribution of the RMS noise. Noise from a total of 9600 (9656) channels was measured for the atmospheric pressure () case, which corresponds to 96% of the total number of OROC pads as described in table 8. We require the waveform baseline to be greater than counts in figure 12. This cut is placed to remove waveforms that have noise cut off at due to the low baseline, which makes the waveform appear to have low noise. This cut removes of the data, and results in 8621 (8645) channels for the atmosphere () case. Correction of the baseline can be applied to individual SAMPA channels, which would mitigate the issue of low baseline and therefore recover these channels. This calibration was not performed during this operation period of TOAD due to time constraints.
@@
The noise distribution is bimodal with a single-bin peak at just above 0 and another peak at . The first peak is caused by FECs that are not properly connected to the OROC. This peak is spatially shown as the grey patches in figure 11. This is demonstrated by measuring the RMS noise from the FECs after they are removed from the OROC and left floating. Figure 13 shows a comparison between the RMS noise for the same subset of 12 channels on one FEC, before and after the card is unplugged from the OROC. When the FEC is floating, all of the measured channels have RMS noise of . This level of noise corresponds to the intrinsic noise level of the SAMPA (), as shown in table 1, which is equivalent to and . A gain of is used for the conversion [10]. For most of the FECs that are well-connected to the OROC, their noise distribution is characterised by the second peak.
@@
The large tail that extends past in figure 12 are shown as distinct noisy regions in figure 11. This is due to a partial connection between the FEC and the OROC which results in a floating ground, as this was not observed in benchtop tests with direct connection to the FECs. The difficulty in connecting the FEC to the OROC, for both the case of improper connection, which left the FEC afloat, and partial connection, which resulted in a floating ground, is due to the legacy OROC connectors used in TOAD. It is observed that the same FEC can gain good performance when moved to a difference slot on the backplane. It is also worth noting that the ALICE ROCs are now end of life, so a future HPgTPC would be designed so as to achieve a more reliable connection to the front-end electronics.
@@
From figure 12, no noticeable difference can be seen in the noise measurements taken in atmosphere (mode 1 of table 7) and 96:4 Ar-CH4 (mode 2). A Kolmogorov-Smirnov (K-S) p-value of is obtained from the unbinned data, which suggests the same underlying distribution.
@@
Additionally, we see that powering of the HV devices does not have a noticeable effect on the noise. This can be seen in figure 14 which compares the noise when the HV devices are off in atmosphere (mode 1 of table 7) and when in 98:2 Ar-CH4 (mode 3). A K-S p-value of is obtained and again suggests the same underlying distribution. As the data for this plot were taken in a different condition and before some noise improvements were made compared to figure 12, no direct comparison could be made. Nevertheless, the bimodal distribution and the positions of the peaks are still the same between the two sets of data. We also carried out a test with the same OROC and gating grid voltages but with the drift cathode powered to (mode 4), and the main features in the noise distribution remained identical.
@@
The contribution of noise from different frequencies can be studied through Discrete Fourier Transformation (DFT) of the waveforms. Figure 15 shows the DFTs from a single channel on a FEC. To avoid random noise, 1000 waveforms are collected and the average of the DFT of 1000 individual waveforms is plotted. Large peaks at and integer multiples of it can be noticed in figure 15. This corresponds to the switching frequency of the FEC DC/DC converter. Less prominently, the smaller peaks at multiples of corresponds to the switching frequency of the FPGA DC/DC converter.
@@
To summarise, the DC/DC components on the PAT boards are major contributors to the noise measured by the FECs. In addition, FECs that were not conductively connected to the OROC would appear to have extremely low noise. High noise channels in the OROC are seen in small regions, usually corresponding to a single FEC, and could be due to improper grounding of the OROC connectors.
@@
To assess whether the level of noise measured in TOAD would be sufficient for a neutrino experiment, we estimate the size of the signal on a FEC that could be generated by a minimally ionising particle. The expected signal-to-noise ratio can be calculated using stopping power , atmospheric argon gas density , and ion pair production energy , giving at atmosphere[18]. At , using a projected gas gain of from [9] and accounting for the change in gas density, one obtains . We take as an approximate track length across an OROC pad, i.e., a channel in a FEC. Simulations have been done to show that the electrons spread across multiple pads, with only 22% lie within of the centre. Furthermore, it takes 7 samples for the charge to be fully deposited. So using from [10], on average, a minimally ionising particle is expected to deposit per channel per sample. The 95 percentile of figure 12 gives a conservative estimate for noise of , which leads to an estimated signal-to-noise ratio of 21.

### v2

@@ 3.2 Firmware design
The PAT card’s main processing unit is a Kintex Ultrascale FPGA package. This device was selected as it provides the required number of serial transceivers to accommodate both readout schemes for ND-GAr whilst satisfying the remaining requirements. The power consumption (based on the AMD power estimate tool) is within the nominal limits for operation in gases at atmospheric pressure or above when actively cooled. Finally, the availability of this package at low cost through CERN and the extensive experience of using it that has been built up at CERN make this FPGA a favourable choice given the cost-driven nature of the design.
In order to take advantage of the FPGA, dedicated firmware had to be developed that allowed data to be reliably and efficiently streamed. Here, we discuss the firmware solution for direct powering, as this was the firmware that was eventually tested with the TOAD test beam. The design has two primary functions. First, the data flow, i.e. how FEC data are reliably received into the software. The second is Control Configuration and Monitoring (CCM) operations. A schematic representation of the firmware design can be seen in figure˜5.
@@ 3.2.1 Data flow
The data stream is then decoded using dedicated decoder firmware based on a Finite State Machine (FSM) that can handle all the different operational modes of the SAMPA. The extracted payload, along with some of the header information (channel number, timestamp, and number of clusters), is converted into a stream that is subsequently sent through a RR based buffering system. The output Ethernet sample additionally contains a single header with some of the header information received from the SAMPA header, as well as the DUNE timestamp for data alignment.
@@
The second option interfaces the SiTCP suite from BeeBeans [15] with IPbus. SiTCP is a lightweight interface that provides streaming capabilities over Transmission Control Protocol (TCP). We developed an interface between the standard pull-request based IPbus firmware and the SiTCP IP core, thereby providing a TCP-based interface for the IPbus suite. While SiTCP includes its own UDP-based interface for CCM, it does not have enough address space for our application.
The third option was used for the power over serial powering mode investigations. The number of serial links on the PAT only allows for a single hardware interface between the in-pressure and out-of-pressure boards for both data and CCM and hence the SiTCP solution can be used in the out-of-pressure electronics. The link between the in-pressure and out-of-pressure electronics still cannot use the SiTCP solution, so it is implemented as an Aurora-based stream interface. In this operation mode, IPbus is not used at all, and the received data are converted into a TCP stream interface using SiTCP.
@@ 3.2.2 Control, Configuration and Monitoring
The firmware includes many registers containing information allowing full telemetry of the hardware from software. The CCM aspect of the firmware’s job is to monitor and set these registers. The design of the PAT cards allows the FPGA to also monitor and change the behaviour of several other components on the board using standardised slow control interfaces (e.g. Inter-Integrated Circuit (I2C), Serial Peripheral Interface (SPI), Power-Management Bus (PMbus)). The list of components can be found in table˜8 along with a short explanation of their function.
@@
In addition to the external hardware components, the CCM part of the firmware implements a variety of standard IPbus registers and memories dedicated to monitoring the data flow or configuring operational modes. A list of these registers can be seen in table˜9.
@@
| Sync. control | Send sync. signal to all FEC sampling windows | 1/FEC group |
@@ 3.3 Software development
The data flow application is based on the DUNE DAQ software suite. For CCM, an IPbus-based framework has been developed, allowing full access to all the in-pressure electronics’ configuration registers. Finally, for timing, the software is based on the framework from the DUNE timing system, which we have wrapped in the bespoke PAT CCM framework since the timing system uses IPbus for configuration as well.
@@ 3.3.1 DUNE DAQ integration
Once the data has been streamed out, it is received by the DUNE DAQ software. While this software already existed prior to this work, significant effort was required to integrate TOAD’s DAQ into the larger DUNE software stack. A schematic of the general data flow within DUNE DAQ can be found in [16]. The integration of the TOAD readout into this framework involved adding the Ethernet packet structure described in section˜3.2.1 into the upstream DAQ. Additionally, the zero-suppression functionality of the FECs allows for the PAT boards to only output data above the threshold we wish to be considered for offline processing. This means that TOAD also demonstrated that the DUNE DAQ software framework worked in untriggered continuous readout mode.
A further difference between TOAD and the other systems using DUNE DAQ is that the output stream data format allows each hit that is streamed out to contain a variable number of ADC samples, while all other systems had fixed numbers of samples per transmitted data packet. This implementation expands the capabilities of the DUNE DAQ software and allows it to handle variable-length data packets [13].
@@ 3.3.2 Control, Configuration and Monitoring software suite
The data from TOAD is received by standard DUNE software [13]. However, the CCM software suite is a new, scalable and modular software framework. The framework has two key base classes to talk to the two classes of objects on the PAT boards: firmware-implemented registers and external devices. The framework is summarised in figure˜8. Each base class contains general methods needed by all registers of that type. Communication to the firmware registers is via IPBus, so the base class contains IPBus-specific methods, while the inherited classes for each specific register contain their own specific bit maps that relate their 32-bits to specific functions. This scheme allows scalability to all electronic devices that use IPBus and similar firmware implementation.
@@
The CCM suite also provides various low-level scripts that act as a Command Line Interface (CLI) for the user to manage the hardware. By combining multiple low-level scripts, one can create high-level scripts that are tailored towards specific use cases. Examples include the noise measurements outlined in section˜4.1 and continuous temperature and power monitoring used in section˜4.2.
@@ 4.1 Noise characterisation
Understanding the noise profile is crucial in verifying that we have a sufficient signal-to-noise ratio to identify the signals we are looking for. We characterise noise as the Root Mean Square (RMS) error of the waveform in the absence of any signal. Prior to measuring the noise and in order to ensure that everything is correctly set up, the PAT cards have been tested for synchronisation. This has been achieved on the table top setup, where the clock skew between the different FEC links has been measured and found to be in the sub-ps regime, indicating that all the FECs were clocked at the same time. The same test with the same outcome has been performed for the trigger signal, which is crucial for collecting noise data.
@@
Noise data were collected by sending trigger signals to the FECs using the high-level script shown in figure˜8. The noise is measured by taking the RMS of a waveform, which corresponds to 500 samples at , for each channel in each FEC. Figure˜9 shows the spatial distribution of the RMS noise of all FECs in the OROC, measured under pressure. Channels with RMS noise less than are coloured grey, while channels with noise above are shown with the highest colour in the scale. The blank regions are caused by FECs that did not respond to any I2C communications.
@@
Figure˜10 plots the distribution of the RMS noise. Noise from a total of 9600 (9656) channels was measured for the atmospheric pressure () case, which corresponds to 96% of the total number of OROC pads as described in table˜1. We require the waveform baseline to be greater than counts in figure˜10. This cut is placed to remove waveforms that have noise cut off at due to the low baseline, which makes the waveform appear to have low noise. This cut removes of the data, and results in 8621 (8645) channels for the atmosphere () case. Correction of the baseline can be applied to individual SAMPA channels, which would mitigate the issue of low baseline and therefore recover these channels. This calibration was not performed during this operation period of TOAD due to time constraints.
@@
The noise distribution is bimodal with a single-bin peak at just above 0 and another peak at . The first peak is caused by FECs that are not properly connected to the OROC. This peak is spatially shown as the grey patches in figure˜9. This is demonstrated by measuring the RMS noise from the FECs after they are removed from the OROC and left floating. Figure˜11 shows a comparison between the RMS noise for the same subset of 12 channels on one FEC, before and after the card is unplugged from the OROC. When the FEC is floating, all of the measured channels have RMS noise of . This level of noise corresponds to the intrinsic noise level of the SAMPA (), as shown in table˜5, which is equivalent to and . A gain of is used for the conversion [11]. For most of the FECs that are well-connected to the OROC, their noise distribution is characterised by the second peak.
@@
The large tail that extends past in figure˜10 is shown as distinct noisy regions in figure˜9. This is due to a partial connection between the FEC and the OROC, which results in a floating ground, as this was not observed in benchtop tests with direct connection to the FECs. The difficulty in connecting the FEC to the OROC, for both the case of improper connection, which left the FEC afloat, and partial connection, which resulted in a floating ground, is due to the legacy OROC connectors used in TOAD. It is observed that the same FEC can gain good performance when moved to a different slot on the backplane. It is also worth noting that the ALICE ROCs are now end of life, so a future HPgTPC would be designed so as to achieve a more reliable connection to the front-end electronics.
@@
From figure˜10, no noticeable difference can be seen in the noise measurements taken in atmosphere (mode 1 of table˜2) and 96:4 Ar-CH4 (mode 2). A Kolmogorov-Smirnov (K-S) p-value of is obtained from the unbinned data, which suggests the same underlying distribution.
@@
Additionally, we see that powering of the HV devices does not have a noticeable effect on the noise. This can be seen in figure˜12 which compares the noise when the HV devices are off in atmosphere (mode 1 of table˜2) and when in 98:2 Ar-CH4 (mode 3). A K-S p-value of is obtained and again suggests the same underlying distribution. The data for this plot were taken in a different condition and before some noise improvements were made to address the known issue of poor grounding in the test beam facility. As such, no direct comparison could be made with figure˜10. Nevertheless, the bimodal distribution and the positions of the peaks are still the same between the two sets of data. We also tested mode 4 (same OROC and gating grid voltages and cathode voltage) for a small subset of the channels, shown in figure˜13. The same data for mode 1 (blue) was used for both figure˜12 and figure˜13, but for the latter figure, only the subset of channels matching those collected for mode 4 were shown. A K-S p-value of 0.76 is obtained and the main features in the noise distribution remained identical.
@@
The contribution of noise from different frequencies can be studied through Discrete Fourier Transformation (DFT) of the waveforms. Figure˜14 shows the DFTs from a single channel on a FEC. To avoid random noise, 1000 waveforms are collected and the average of the DFT of 1000 individual waveforms is plotted. Large peaks at and integer multiples of it can be noticed in figure˜14. This corresponds to the switching frequency of the FEC DC/DC converter. Less prominently, the smaller peaks at multiples of correspond to the switching frequency of the main DC/DC converter.
@@
To summarise, the DC/DC components on the PAT boards are major contributors to the noise measured by the FECs. However, since this iteration of the board is not the one expected to be deployed in ND-GAr in a future revision, more care will be taken with regard to the choice of the DC/DC converter to not limit the physics capability of the detector. In addition, FECs that were not conductively connected to the OROC would appear to have extremely low noise. High noise channels in the OROC are seen in small regions, usually corresponding to a single FEC, and could be due to improper grounding of the OROC connectors.
@@
To assess whether the level of noise measured in TOAD would be sufficient for a neutrino experiment, we estimate the size of the signal on a FEC that could be generated by a minimally ionising particle. The expected signal-to-noise ratio can be calculated using stopping power , atmospheric argon gas density , and ion pair production energy , giving at atmosphere [17]. At , using a projected gas gain of from [7] and accounting for the change in gas density, one obtains . We take as an approximate track length across an OROC pad, i.e., a channel in a FEC. Simulations have been done to show that the electrons spread across multiple pads, with only 22% lie within of the centre. Furthermore, it takes 7 samples for the charge to be fully deposited. So using from [11], on average, a minimally ionising particle is expected to deposit per channel per sample. The 95 percentile of figure˜10 gives a conservative estimate for noise of , which leads to an estimated signal-to-noise ratio of 21.